# Temporally Consistent Factuality Probing for Large Language Models

### v1

@@ A.2.6 GPT-4 Evaluation Results
Due to a minimal sample set, we are short of reporting either the temporally-consistent-factuality or temporal-consistency for this test. Metrics like temporal consistency and temporally consistent factuality require a sufficient number of pair of samples for a subject-relation to calculate those. The other factors are temporal direction and entity pairs, which are to be considered while calculating. The 100 individual samples might be enough to gauge the preliminary assessment of temporal factuality but statistically insufficient for reporting temporally consistent factuality. We leave to improve the temporally-consistent-factuality of GPT-4 for future research via methods that may not require gradient updates. Due to a minimal sample set, we are short of reporting either the temporally-consistent-factuality or temporal-consistency for this test. Metrics like temporal consistency and temporally consistent factuality require a sufficient number of pair of samples for a subject-relation to calculate those. The other factors are temporal direction and entity pairs, which are to be considered while calculating. The 100 individual samples might be enough to gauge the preliminary assessment of temporal factuality but statistically insufficient for reporting temporally consistent factuality. We leave to improve the temporally-consistent-factuality of GPT-4 for future research via methods that may not require gradient updates.
S.No. Prompt Correct Answer Model’s Response
1 Black panther won the best costume design academy award succeeded by little women little women
2 The FIFA u-17 world cup was hosted by canada immediately after China Scotland
@@
A.3 Hyperparameters

### v2

@@ A.4.2 Claude-3 Evaluation Results
Due to a minimal sample set, we are short of reporting either the temporally-consistent-factuality or temporal-consistency for this test. Metrics like temporal consistency and temporally consistent factuality require a sufficient number of pair of samples for a subject-relation to calculate those. The other factors are temporal direction and entity pairs, which are to be considered while calculating. The 100 individual samples might be enough to gauge the preliminary assessment of temporal factuality but statistically insufficient for reporting temporally consistent factuality.
Model Settings Temp-fact
GPT-4 GPT-4-0125-preview 36.00
Claude-3 Claude-3-Opus-20240229 20.00
Table 10: Temporal factuality (in %) of two commercial LLMs, GPT-4 and Claude-3 through the minimal sample set.
S.No. Prompt Correct Answer GPT-4 Response
1 Android Gingerbread was released by google immediately after android Froyo Froyo
2 Indian band Euphoria released album Sharnagat right after album Item MeHFuz
@@
A.4.3 CoTSeLF’s Applicability
Applicability across commercial LLMs. Commercial model’s such as GPT-4 and Claude-3’s temporal-factuality remains suboptimal, thus necessitating strategies like CoTSeLF or similar enhancements to augment its temporal-factuality. CoTSeLF advances an RL-based fine-tuning methodology that refines the model’s parameters via a custom-made fine-tuning procedure. Almost universally, advancements in parameter fine-tuning methodologies are contingent upon access to the model’s architecture. GPT-4 and Claude-3, by restricting direct architectural access, necessitate that such methodologies could demonstrate their effectiveness only on open-source models. Thus, we could not assess CoTSeLF’s efficacy on GPT-4 or Claude-3 models due to the inaccessibility of its architecture for open fine-tuning.
Considering that GPT-4 adhered to a decoder-based transformer architecture (up to GPT-2) before the later versions transitioned to a commercial model, there is an intense anticipation that CoTSeLF, particularly CTSRL, will serve as an efficacious approach to augmenting the closed-source model’s capabilities in improving temporally-consistent-factuality as in case of various versions of LLaMA model under experimentation.
An alternative approach involves examining methods that do not necessitate access to the model’s architecture for parameter updates, acknowledging that typically, strategies not involving parameter modifications exhibit suboptimal performance compared to those that do, like CoTSeLF. Future research should focus on enhancing the temporally-consistent factuality of commercial LLMs through techniques that might bypass the need for gradient updates.
Applicability across domains. The establishment of consistent temporal reasoning capabilities is poised to profoundly influence their applicability across a spectrum of complex fields, including healthcare and the legal sector. Specifically, within the legal domain, the temporally reliant nature of case precedents and statutory amendments demands the reliable and precise retrieval of information. LLMs must avoid offering divergent interpretations of precedents for specific case types to prevent severe repercussions and undermine trust in the judicial system. The precedence of rules, articles, and cases, being temporally bound, demands their consistent retrieval when employing LLMs in legal settings. Similarly, in healthcare, the prognostication of diseases is time-sensitive, mandating consistent and precise forecasts for diagnosing and formulating treatment plans, leading to enhanced trust in the usage of AI-based medical solutions. Techniques like CoTSeLF could be tailored to meet such requirements.
A.5 Hyperparameters